A student wearing frictionless roller skates moving at 1.2 m/s on a horizontal surface is then pushed by a friend 6.54 m with a constant force of 45 N. If her final kinetic energy is 352 J, what is the mass of the student? 



1
Expert's answer
2021-06-07T08:32:08-0400

"KE=mv^2/2\to m=2\cdot KE/v^2"


"s=\frac{v^2-v_0^2}{2a}\to v^2=2as+v_0^2=2\cdot \frac{F}{m}\cdot s+v_0^2"


"m=\frac{2\cdot KE}{2\cdot \frac{F}{m}\cdot s+v_0^2}\to m=\frac{2KE-2Fs}{v_0^2}="


"=\frac{2\cdot 352-2\cdot 45\cdot 6.54}{1.2^2}=80.14\ (kg)" . Answer